### v1

# Temperature overshooting in the Mpemba effect of frictional active matter

Abstract
The traditional Mpemba effect refers to an anomalous cooling phenomenon when an initial hotter system cools down faster than an initial warm system. Such counterintuitive behavior has been confirmed and explored across phase transitions in condensed matter systems and also for colloidal particles exposed to a double-well potential. Here we predict a frictional Mpemba effect for a macroscopic body moving actively on a surface governed by Coulomb (dry) friction. For an initial high temperature, relaxation towards a cold state occurs much faster than that for an intermediate initial temperature, due to a large temperature overshooting in the latter case. This frictional Mpemba effect can be exploited to steer the motion of robots and granules.
The Mpemba effect – first reported in antiquity in Aristotle’s Metheorologica [1] and later famously associated with Erasto Mpemba himself, who observed it while freezing ice cream [2] – refers to the counterintuitive phenomenon in which a system initially at a higher temperature reaches its steady state faster than an identical system starting from a lower temperature when both are quenched into the same cold environment. Controversial by nature, this effect has attracted broad attention and has been observed since then in a wide range of systems. Beyond the canonical case of water freezing, notable instances include but are not limited to colloidal suspensions [3, 4, 5] and quantum systems [6, 7, 8, 9]. Nevertheless, despite its consistent phenomenological signature, the underlying mechanisms appear to be system-specific and remain the subject of ongoing debate [10], lacking even a universal heuristic – let alone a theoretical – explanation.
One of the leading theoretical explanations for the Mpemba effect invokes a complex energy landscape featuring multiple metastable states [11, 12]. For simplicity, such landscapes are often visualized as effective potentials with just two minima – a local and a global one [3, 10]. In this framework, a system initialized at a higher temperature is less likely to be found near the local minimum, and thus less likely to become transiently trapped there. Conversely, a system starting from a lower temperature is more probable to occupy a local minimum, resulting in a delayed relaxation to the global one.
While this explanation provides an intuitive understanding of Mpemba-like behavior, it must be adapted for systems where no well-defined energy landscape exists, such as those governed by non-conservative forces or subject to strong nonequilibrium driving. A prominent class of such systems is active matter, characterized by the persistent injection of energy from the environment or internal sources into directed (self-propelled) motion [13, 14]. In these cases, even the notion of temperature becomes ambiguous, and a fundamental question arises: how [15] should temperature be defined in systems far from equilibrium?
In this study, we report the observation of the Mpemba effect in active systems with dry friction, such as active granular matter composed of vibrating robots [16, 17, 18, 19], where particles experience solid-solid contact with the frictional surface (Fig. 1(a)). In contrast to colloidal systems – including active colloids [20] – where the Mpemba effect emerges under velocity-dependent Stokes friction and confinement by an external potential, the system studied here evolves without any external potential and is instead subject to velocity-independent Coulomb, or dry friction [21, 22]. Therefore, we refer to it as a frictional Mpemba effect. The intrinsic interplay between dry friction and active self-propulsion gives rise to several distinct behaviors, observed both theoretically and experimentally, including a continuous switching between diffusive and accelerated motion [23] and a frictional-induced phase separation mechanism [24]. Notably, the steady-state velocity distribution in such system deviates significantly from the Gaussian (Maxwell-Boltzmann) form, which is often considered a prerequisite for observing the Mpemba effect, e.g. in granular fluids with non-elastic collisions [25]. The Mpemba effect observed in our study arises due to an overshooting phenomenon, in which the system temporarily reaches the temperatures below that of the steady state. This overshooting has been identified in other systems as a possible key mechanism underlying anomalous thermal relaxation that facilitates the occurrence of Mpemba-like behavior [12, 26].
Figure 1: (a) An illustrative example of active granular particle (vibrating robot), subjected to dry friction due to the solid-solid contact between the surface and its legs. (b)-(d) Trajectories sketched to illustrate different temperatures. The motion occurs due to an interplay between active force and dry friction force, while the temperature regulates the impact of the fluctuations, resulting in the gradual trajectories for cold temperatures (b) and in fluctuating trajectories when increasing the temperature for warm (b) and hot (c). (e) Dependence of the system temperature $T$ on the noise parameter $\epsilon$. Colored rectangle projections mark the cold, warm, and hot states, highlighting the associated ranges of temperature $T$ and noise strength $\epsilon$.
Here we consider a minimal model to study an inertial active particle with mass $m$ subject to a dry friction force $\sigma(v)=\Delta_{C}\operatorname*{sign}(v)$, where $\Delta_{C}$ is the dry friction coefficient. The dynamics is described by a one-dimensional Langevin equation for the particle velocity $v(t)=\dot{x}(t)$,
$m\dot{v}(t)=-\sigma(v(t))+\sqrt{2K}\xi(t)+n(t)f,$ (1)
where $t$ is time, $\xi(t)$ is a Gaussian white noise process with zero mean and correlations $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t^{\prime}-t)$, $K$ is the white noise strength, the time-dependent term $n(t)f$ is the active force with activity amplitude $f$ and the stochastic process $n(t)$ chosen as an Ornstein-Uhlenbeck process:
$\dot{n}(t)=-\frac{n(t)}{\tau}+\sqrt{\frac{2}{\tau}}\eta(t).$ (2)
Here $\tau$ is the persistence time, and $\eta(t)$ represents Gaussian white noise with zero mean and correlations $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t^{\prime}-t)$. This choice of the active force is therefore typically referred to in literature as active Ornstein-Uhlenbeck particle dynamics [27, 28, 29, 30, 31]. In this study, we focus on the overall impact of stochastic contributions, both white and Ornstein-Uhlenbeck noise. Since these noise sources can stem from related physical mechanisms [16], it is natural to regulate the temperature through a single composite parameter that reflects their combined effect.
In what follows, we set $\Delta_{C}$, $\sqrt{\tau K}/\Delta_{C}$, $\tau K/m\Delta_{C}$ as units of force, time, and length, respectively. Under this rescaling, the behavior of the system reads:
$\dot{v}(t)$ $=$ $-\operatorname*{sign}(v(t))+n(t)f_{0}+\sqrt{2\epsilon}\,\xi(t),$ (3a)
$\dot{n}(t)$ $=$ $-\epsilon{n(t)}+\sqrt{2\epsilon}\,\eta(t).$ (3b)
It depends solely on two key dimensionless quantities: the reduced activity amplitude $f_{0}=f/\Delta_{C}$, which quantifies the strength of the active force relative to dry friction and is fixed at $f_{0}=0.5$ throughout this study, and the noise parameter $\epsilon=\sqrt{K/\tau}/\Delta_{C}$ which controls both sources of noise in the system – the white noise with strength $K$ and the active noise (Ornstein-Uhlenbeck) characterized by the persistence time $\tau$. Through its presence in the stochastic contributions of Eq. (3), $\epsilon$ thus governs the overall impact of fluctuations on the particle dynamics and thus determines the temperature of the system. In what follows, we adopt the kinetic temperature, defined as
$T=\frac{1}{2}\frac{m\langle v^{2}\rangle}{k_{B}},$ (4)
as our measure of the system’s temperature.
In the low temperature regime sketched in Fig. 1(b), the dynamics are gradual, while increasing the temperature in Fig. 1(c)–(d) amplifies the role of fluctuations, resulting in progressively noisier and more entangled trajectories. In Fig. 1(e), we demonstrate how temperature can be controlled by the noise parameter $\epsilon$.
Figure 2: (a) Steady-state velocity probability distribution $P_{\rm st}(v)$ for various noise parameters $\epsilon$ and fixed activity amplitude $f_{0}=0.5$. (b) Kurtosis of the steady-state distribution as a function of noise parameter $\epsilon$ at a fixed activity amplitude $f_{0}=0.5$. Color phases correspond to those depicted in Fig. 1(e), and the dashed vertical lines correspond to the noise parameters of density profiles shown in panel (a).
In case of dry friction, the particle does not move until the force acting on it exceeds the threshold value (which is $\Delta_{C}$, or 1 in our dimensionless units) of the friction force. Thus, the particle’s motion can be initiated by either Brownian noise or an active force. We note that the active force contribution to the particle motion is unaffected by temperature, as the steady-state distribution of the stochastic process $n$, $P_{\rm st}(n)=e^{-n^{2}/2}/\sqrt{2\pi}$ is independent of $\epsilon$. However, the probability that the active force exceeds the threshold,
[…]
remains small – but non-negligible – due to our specific choice of the activity amplitude $f_{0}=0.5$.
At high noise parameter $\epsilon$, the motion is predominantly driven by Brownian noise as it dominates the activity effects, leading to a characteristic Laplacian (V-shaped) steady-state velocity distribution $P_{\rm st}(v;\epsilon)$ in log scale (Fig. 2(a)), characteristic for Brownian particles subjected to dry friction [21, 32]. In contrast, at low $\epsilon$ the Brownian noise is suppressed by dry friction, and motion is only possible due to rare yet contributing fluctuations in the active force leading to accelerated motion. Such periods of accelerated motion, albeit rare, are lasting much longer due to the higher persistence of the Ornstein-Uhlenbeck process (3b) at low noise parameter $\epsilon$, therefore leading to heavy distribution tails and a sharp peak at zero in the steady-state velocity distribution $P_{\rm st}(v;\epsilon)$ (Fig. 2(a)). The kurtosis $\kappa=\langle v^{4}\rangle/(\langle v^{2}\rangle)^{2}-3$, shown in Fig. 2(b), illustrates the transition from activity-dominated dynamics at low $\epsilon$ to noise-dominated behavior at high $\epsilon$. As noise increases, $\kappa$ decreases from large values, reflecting heavy-tailed velocity distributions, and approaches the kurtosis of a Laplacian distribution $\kappa_{\textrm{\tiny{L}}}\!=\!3$. This is intrinsically different from most previously reported Mpemba-like effects caused by non-Gaussian velocity distributions, which occurred in the low-kurtosis regime [25, 33].
Figure 3: (a) Relaxation dynamics of the distance measure $\mathcal{D}(t)$. (b) Relaxation time in the temperature range $T_{\rm init}/T_{\rm bath}\in[1,900]$. The blue rectangle indicates the temperature range where the Mpemba effect (a reduction in relaxation time) is observed. (c) Time evolution of system temperatures for warm and hot initial temperatures considered in panel (a).
[…]
and the relaxation time $t_{\rm relax}$ is defined as the moment when the distance reaches $\mathcal{D}(t_{\rm relax})=\mathcal{D}_{\rm relax}=2\cdot 10^{-3}$, comparable to the noise level.
The distance evolution shown in Fig. 3(a) demonstrates a strong Mpemba effect, since the initially hotter system reaches the steady state first despite starting at a greater distance from the bath distribution than the initially colder systems. The decrease in cooling time with increasing initial bath temperature, shown in Fig. 3(b), reveals the temperature range in which the Mpemba effect occurs.
But what are the factors that give rise to the Mpemba effect? To gain a better understanding of this phenomenon, we track the time evolution of the system (kinetic) temperature. At the first stages of the relaxation process depicted in Fig. 3(c), the system is being slowed down (i.e., cooled in terms of temperature) by the dry friction. In contrast to velocity-dependent types of friction such as fluid friction, dry friction does not depend on velocity. Therefore, the efficiency of cooling does not depend on the system temperature, unlike in the case of fluid friction, where faster-moving objects are slowed down (cooled) more quickly than slower ones. Since in the considered bath regime, the white noise contribution is negligible and activity enters only later due to the changed protocol (with the mean activation time $t_{\times}\approx 1/\mathcal{P}(nf_{0}\geq 1)$ indicated by the dashed vertical line in Fig. 3(c)), the initially colder system overshoots the bath temperature (shown as a dashed horizontal line in Fig. 3(c)). As a result, it extends the time period required to relax to the bath temperature. This overshooting effect is less pronounced when starting from high initial temperatures, as the corresponding fat tails in the velocity distribution shown in Fig. 2(a) – characteristic of low and high temperatures but absent at intermediate temperatures between them – enable a more direct relaxation pathway toward the target distribution.
In conclusion, the observed Mpemba effect emerges as a generic consequence of nonlinear friction dynamics, which inherently lead to the temperature overshooting during relaxation. Our results highlight the fundamental role of the interplay between friction and activity, offering new perspectives on anomalous cooling phenomena in this class of systems. The simplicity and generality of our model suggest that the Mpemba effect extends beyond conventional temperature definitions and can manifest itself in broader classes of observables which are characterized by fat tails in their probability distributions. This opens new avenues for exploring generalized Mpemba-like behavior in various physical settings, including but not limited to active particles moving through disordered environments [34], relaxation processes in dense granular flows [35] or in controlled protocols such as thermodynamically equidistant temperature quenches [36, 37]. In contrast to other Mpemba-like effects observed previously in granular systems [25, 33, 38, 39, 40], the frictional Mpemba effect found here arises due to the overshooting of highly mobile particles as they relax toward the steady state. Thus, the phenomenon is amenable to experimental observation at the macroscopic scale and is of great practical significance for controlling dynamical properties across a wide range of active platforms, such as vibrobots on frictional surfaces or swarm robotics operating in complex environments [41].
Acknowledgment. HL acknowledge the financial support by Deutsche Forschungsgemeinschaft (German Research Foundation), Project LO 418/25-1.